A 12 pole, 3-phase alternator drives at a speed of 500 r.p.m supplies power to an 8 pole, three-phase induction motor. If the slip of the motor at full load speed is 4 %, calculate the full load of the motor. 


1
Expert's answer
2021-08-30T01:54:43-0400

We can determine this as follows:


Given that, speed of alternator (NS) = 500

Number of poles (P) = 12

\( \Rightarrow 500 = \frac{{120f}}{12} \Rightarrow f = 50\;Hz\)

The number of poles in induction motor = 8

The synchronous speed of the induction motor is,

\({N_s} = \frac{{120 \times 50}}{8} = 750\;rpm\)

Slip (s) = 4 % = 0.04

Full load speed, Nr = 750(1 – 0.04) = 727.5 rpm

Rounded off, we get 728 rpm